1.y(3x^3-x+y)dx+x^2(1-x^2)dy=0
2.y(2x+y^2)dx+x(y^2-x)dy=0
1
Expert's answer
2021-07-22T12:06:58-0400

1.

Solution;

Rewrite the solution as;

"\frac{dy}{dx}x^2(1-x^2)+(3yx^3-xy+y^2)=0"

"\frac{dy}{dx}x^2(1-x^2)+y(3x^3-x)=-y^2"

Divide by "x^2(1-x^2)"

"\frac{dy}{dx}+\frac{y(3x^3-x}{x^2(1-x^2}=\frac{-y^2}{x^2(1-x^4}"

Divide by -y2

"-\frac{dy}{dx}+\frac{(1-3x^2)}{(x-x^3)y}=\frac{1}{x^2-x^4}"

Let v="\frac 1y" so that "\frac{dv}{dx}=\frac{-dy}{dxy^2}"

Substitute in the equation;

"\frac{dv}{dx}+\frac{(1-3x^2)v}{x-x^3}" ="\frac{1}{x^2-x^4}"

Find the integrating factor;

I.F="e^{\int{\frac{1-3x^2}{x-x^3}}}=e^{ln(x^3-x)}=x^3-x"

Multiply all through with the integrating factor;

Substitute 3x2-1="\frac{d(x^3-x)}{dx}"

"(x^3-x)\frac{dv}{dx}+\frac{d(x^3-x)v}{dx}=\frac{x^3-x}{x^2-x^4}"

Apply reverse product rule:"\frac{f(dg)}{dx}+\frac{g(df)}{dx}=\frac{d(fg)}{dx}"

We have;

"\frac{d(x^3-x)v}{dx}=\frac{x^3-x}{x^2-x^4}"

Integrate both sides with respect to x;

("x^3-x)v=-ln(x)+c"

"v=\frac{-ln(x)+C}{x^3-x}"

But "v=\frac1y"

"\frac1y=\frac{ln(x)+C}{x-x^3}"

Simplify;

y="\frac{x-x^3}{ln(x)+C}"

2.

Solution;

By distribution,the equation may be written as;

"(2xy+y^3)dx+(xy^2-x^2)dy=0"

Check if the equation is exact,

M="2xy+y^3 ;\frac{dM}{dy}=2x+3y^2"

N="xy^2-x^2 ;\frac{dN}{dx}=y^2-2x"

"\frac{dM}{dy}\neq\frac{dN}{dx}" ,the equation is not exact.

Mx-Ny=2x2y+y3x-xy3+x2y=3x2y

Take an integrating factor given as;

I.F="\frac{1}{Mx-Ny}=\frac{1}{3x^2y}"

Multiply the given equation with the integrating factor ;

"\frac{2x+y^2}{3x^2}dx+\frac{y^2-x}{3xy}dy=0"

The general solution of the problem will be given by;

"\int{Mdx}+\int"(Terms of N which are independent of x)

"\int{\frac{2x+y^2}{3x^2}}dx+\int{\frac{-1}{3y}}dy=C"

"\frac{2ln(x)}{3}-\frac{y^2}{3x}-\frac{ln(y)}{3}=C"